You are ordering softballs for 2 softball leagues. The size of a softball is measured by its circumference. The pony league uses an 11-inch softball priced at $3.50. The junior league uses a 12-inch softball priced at $4.00. The bill smeared in the rain but you know the total was 80 softballs for $305.How many of each size did you order?
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

P+J=80 OR P=80-J
3.5P+4J=305
3.5(80-J)+4J=305
280-3.5J+4J=305
.5J=305-280
.5J=25
J=25/.5
J=50 NUMBER OF JUNIOR BALLS BOUGHT.
80-50=30 NUMBER OF PONY BALLS BOUGHT.
PROOF:
3.5*30+4*50=305
105+200=305
305=305

Answer by stanbon(48568)

The pony league uses an 11-inch softball priced at $3.50. The junior league uses a 12-inch softball priced at $4.00. The bill smeared in the rain but you know the total was 80 softballs for $305.How many of each size did you order?
---------------
Let # of 11-in balls be "x" ; value = 3.5x dollars
-----------------
# of 12-in balls is "80-x" ; value = 4(80-x)=320-4x dollars
-----------------
EQUATION:
3.5x +320-4x = 305
-0.5x = -15
x = 30 (# of 11-in. balls)
80-x = 50 (# of 12-in balls)
=============================
Cheers,
Stan H.